OpenStudy (anonymous):

find the indicated function composition and evaluate. find (g*f)(2) f(x)=squareroot of (2x+9) g(x)=4x^2

zepdrix (zepdrix):

Hey glea :) Is that a multiplication star? Or did you mean for it to be the composition symbol? \(\Large\rm (g\circ f)(2)\)

OpenStudy (anonymous):

what you typed dont know how to make one

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{f(x)=\sqrt{2x+9}},\qquad\qquad g(x)=4x^2\]

zepdrix (zepdrix):

I prefer to use this notation, I think it makes more sense. \[\Large\rm (g \circ f)(x)=\quad g(f(x))\]Hopefully you've seen it before :o

OpenStudy (anonymous):

ya ive seen it both ways

zepdrix (zepdrix):

So we're taking our function g,\[\Large\rm g(\color{royalblue}{x})=4(\color{royalblue}{x})^2\]And instead of plugging x into the function, we're plugging f(x) into it,\[\Large\rm g(\color{orangered}{f(x)})=4(\color{orangered}{f(x)})^2\]

zepdrix (zepdrix):

\[\Large\rm g(\color{orangered}{f(x)})=4(\color{orangered}{\sqrt{2x+9}})^2\]

zepdrix (zepdrix):

And then we're evaluating this composition at x=2,\[\Large\rm g(\color{orangered}{f(2)})=4(\color{orangered}{\sqrt{2\cdot 2+9}})^2\]

zepdrix (zepdrix):

What'd you think? :d Any of that too crazy?

OpenStudy (anonymous):

so 52?

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

and you do the same by plugging g into f tosee if they come out the same?

zepdrix (zepdrix):

Mmmmm no. Unless you're trying to verify that g(x) and f(x) are inverse functions. Were we trying to do something like that? :o Otherwise we're done with the problem.

OpenStudy (anonymous):

no didnt know if i had to show f(g(2))

zepdrix (zepdrix):

We solved this: \(\Large\rm (g\circ f)(2)\) That would correspond to a different problem: \(\Large\rm (f\circ g)(2)\) So nah, we don't need to worry about that :)

OpenStudy (anonymous):

k thanks